# Torsion of Rational Elliptic Curves over the Cyclotomic Extensions of $\mathbb{Q}$

### v5

@@ 5. Relaxing the Conditions
In summary, the cases $p=2$ and $p=3$ are well-covered by Mazur’s theorem and Najman’s work, respectively, while the cases $p=5,7,11$ are covered by Gužvic and Vukorepa. Our approach for $p>5$ provides a more general classification, along with methods to further refine the possibilities for specific primes.
@@ 6. Realization of Torsion Subgroups
Let $E/\mathbb{Q}$ be an elliptic curve. Let $p>3$ be a prime such that $p-1$ is not divisible by $3,4,5,7,11$. If $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is isomorphic to one of the following groups $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z},\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/12\mathbb{Z}$ or $\mathbb{Z}/16\mathbb{Z}$, then $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}=E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}$.
@@
Assume that $E$ is in short Weierstrass form. Then $\mathbb{Q}(E[2])$ is the splitting field of the cubic $x^{3}+Ax+B$. Since $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ has full $2$-torsion, the field $\mathbb{Q}(E[2])$ must be a subfield of $\mathbb{Q}(\zeta_{p})$. Then the cubic cannot be irreducible, since $p-1$ is not divisible by $3$. This means that either the cubic has three rational roots, or it has one rational root and two other roots from $\mathbb{Q}(\sqrt{-p})$. This implies that $E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}$ also has full $2$-torsion. Ultimately, $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}=E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}$.
Secondly, let $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\cong\mathbb{Z}/16\mathbb{Z}$. Let $P=(x_{0},y_{0})\in E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ be a point of order $16$. Then, the extension degree of $\mathbb{Q}(x_{0})$ is equal to the number of elements in the set $\{x_{0}^{\sigma}:\sigma\in\operatorname{Gal}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})\}$, and it also divides $n$. In total, there are exactly $8$ points of order $16$ in $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$, and their $x$-coordinates form a set of $4$ elements. This means that $\{x_{0}^{\sigma}:\sigma\in\operatorname{Gal}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})\}$ has at most $4$ elements. It cannot have $3$ or $4$ elements because we assumed $n$ is not divisible by them. This implies that $\mathbb{Q}(x_{0})$ is at most a degree $2$ extension, so $x_{0}\in\mathbb{Q}(\sqrt{-p})$. Also, $[\mathbb{Q}(x_{0},y_{0}):\mathbb{Q}(x_{0})]\leq 2$, so the extension degree of $\mathbb{Q}(x_{0},y_{0})$ is a divisor of $4$, but it also divides $n$, so it is at most a degree $2$ extension. This implies $P\in E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}$. Since $P$ generates $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$, we obtain $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}=E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}$. ∎
This theorem shows that if $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is not one of the groups in Mazur’s theorem, then we should consider only the quadratic extension and look for torsion points there. These three groups appear in Theorem 2.2, so we know that they are realizable over infinitely many elliptic curves. However, we are interested in whether they are realizable over the quadratic fields $\mathbb{Q}(\sqrt{-p})$ for primes $p$ such that $p-1$ is not divisible by $3$, $4$, $5$, $7$, or $11$.
@@
Torsion of elliptic curves over $\mathbb{Q}(\sqrt{d})$ is studied by Trbović in [25] for $1000<d<$ 100. If a similar study is conducted in the future for $\mathbb{Q}(\sqrt{-d})$ instead, this would lead to a classification of torsion over $\mathbb{Q}(\zeta_{p})$ for small primes like $p=47$ or $p=59$, where $p-1$ is not divisible by $3,4,5,7,11$.
The realization of $\mathbb{Z}/16\mathbb{Z}$ as a torsion subgroup over $\mathbb{Q}(\zeta_{p})$ remains an open problem in general. In order for this group to appear as a torsion subgroup, we need to find a prime $p$ such that $E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}\cong\mathbb{Z}/16\mathbb{Z}$ when $p\equiv 3\pmod{4}$. According to Theorem 1.3 in Banwait and Derickx [1], such a realization does not occur for any prime $p<800$ with $p\equiv 3\pmod{4}$. This rules out all the primes considered in Theorem 1.1.
Their result further shows that among the primes $p\equiv 1\pmod{4}$, the only prime less than $800$ for which this torsion structure is realized is $p=41$. For example, the elliptic curve 266910.ck6 satisfies $E(\mathbb{Q}(\sqrt{41}))_{\text{tors}}\cong\mathbb{Z}/16\mathbb{Z}$ (available on LMFDB). This fact, combined with Theorem 2.3, implies that the torsion subgroup $E(\mathbb{Q}(\zeta_{41}))_{\text{tors}}$ is isomorphic to $\mathbb{Z}/16\mathbb{Z}$.
Now, we turn our attention to the subgroups that arise when we relax the conditions in Theorem 1.1 and are realized over only finitely many elliptic curves. These groups are of the form $\mathbb{Z}/N\mathbb{Z}$, where the modular curve $X_{0}(N)$ has genus greater than zero and possesses non-cuspidal rational points. For example, we are not interested in the case $N=20$, even though $X_{0}(20)$ has genus one and six rational points, because all of these points are cusps. Therefore, there is no rational elliptic curve with a $20$-isogeny.
@@
where $A,B$ are rationals and $d$ is a square-free integer. This implies that the field of definition $\mathbb{Q}(P_{d})$ is contained in $\mathbb{Q}(P,\sqrt{d})$. As an example, we have the elliptic curve 121.a2 and its $11$-torsion is defined over $\mathbb{Q}(\zeta_{11})$. If we consider its $-11$-quadratic twist 121.c1, then its $11$-torsion is defined over $\mathbb{Q}(\zeta_{11})^{+}$. If we consider its $-3$-quadratic twist 1089.i2, then its $11$-torsion is defined over $\mathbb{Q}(\zeta_{33})^{+}$. This data is available on LMFDB.
@@
In the cases $N=67$ and $N=163$, the LMFDB database does not specify the exact field of definition of the $N$-torsion. However, for $N=67$, it does state that the extension degree of the field is equal to $(N-1)/2$. Combining this fact with the discussion in the proof of Lemma 4.7, we can conclude that the field of definition is $\mathbb{Q}(\zeta_{N})^{+}$, as it is the unique subfield of $\mathbb{Q}(\zeta_{N})$ of degree $(N-1)/2$ over $\mathbb{Q}$. For $N=163$, all we can say is that the field of definition is contained in $\mathbb{Q}(\zeta_{N})$.
@@
If $j(E)=-7\cdot 137^{3}\cdot 2083^{3}$, the situation becomes more intriguing. We know that the field of definition of the $37$-torsion is an abelian number field contained in $\mathbb{Q}(\zeta_{n})$ for some positive integer $n$. To determine this minimal value of $n$, we apply Corollary 4.6, using the conductor of the minimal quadratic twist available on the LMFDB. The database also indicates that the extension degree of the field of definition is $36$ over $\mathbb{Q}$. From this, we deduce that $n$ must divide $1295$.
To confirm that $n$ cannot be smaller, we use SageMath to factor the $37$-division polynomial of $E$ over $\mathbb{Q}$. This polynomial factors into four irreducible components over $\mathbb{Q}$: one of degree $18$ and three of degree $222$. Let $P=(x_{0},y_{0})\in E(\overline{\mathbb{Q}})_{\text{tors}}$ be the generator of the kernel of the $37$-isogeny from $E$ to another elliptic curve. Then the degree $[\mathbb{Q}(x_{0}):\mathbb{Q}]$ must divide $36$. Hence, $\mathbb{Q}(x_{0})$ has degree $18$, and its minimal polynomial corresponds to the first factor of the $37$-division polynomial.
We then examine the splitting field of this polynomial and observe that $5$, $7$, and $37$ are ramified in the extension. This confirms that the minimal $n$ for which the field of definition is contained in $\mathbb{Q}(\zeta_{n})$ is indeed $1295$ for the minimal quadratic twist corresponding to this $j$-invariant.
@@
14 $-3^{3}\cdot 5^{3}$ $-7$ 49.a2 $7^{2}$ $\mathbb{Q}(\zeta_{7})^{+}$
$3^{3}\cdot 5^{3}\cdot 17^{3}$ $-7$ 49.a1 $7^{2}$ $\mathbb{Q}(\zeta_{7})^{+}$
@@
$23^{3}\cdot 5^{3}/$ 2 Minimal 162.b4 $2\cdot 3^{4}$ $\mathbb{Q}(\zeta_{9})$
@@
We have already discussed $\mathbb{Z}/16\mathbb{Z}$ earlier in this chapter and provided an example where $16$-torsion is realized over $\mathbb{Q}(\zeta_{41})$. For the other cases, the elliptic curve 147.c2 satisfies $E(\mathbb{Q}(\zeta_{7})^{+})_{\text{tors}}\cong\mathbb{Z}/13\mathbb{Z}$, the elliptic curve 14.a4 satisfies $E(\mathbb{Q}(\zeta_{7})^{+})_{\text{tors}}\cong\mathbb{Z}/18\mathbb{Z}$, and the elliptic curve 11.a3 satisfies $E(\mathbb{Q}(\zeta_{11})^{+})_{\text{tors}}\cong\mathbb{Z}/25\mathbb{Z}$.

### v6

@@ 5. Relaxing the Conditions
In summary, the cases $p=2$ and $p=3$ are well-covered by Mazur’s theorem and Najman’s work, respectively, while the cases $p=5,7,11$ are covered by Gužvić and Vukorepa. Our approach for $p>3$ provides a more general classification, along with methods to further refine the possibilities for specific primes.
We conclude this section by presenting a comprehensive table that summarizes Theorems 1.1, 5.1, 5.2, and 5.3. This table includes the conditions on the primes, the exceptional torsion subgroups (i.e., those not appearing in Mazur’s theorem) that can occur, and specifies for which groups explicit examples are known or remain unknown. While some examples have been provided within the proofs of these theorems, a more detailed discussion of these examples, along with additional constructions, will be presented in the next section.
Table 1. Summary of exceptional torsion subgroups over cyclotomic fields and their realizability
Divisibility conditions on primes $p>3$ Possible exceptional groups Groups without known examples
$p-1$ is not divisible by 333, 444, or 555 $\mathbb{Z}/16\mathbb{Z}$, $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z}$, $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/12\mathbb{Z}$ $\mathbb{Z}/16\mathbb{Z}$
$p-1$ is not divisible by 333 or 444 $\mathbb{Z}/N\mathbb{Z}$ with $N=11,16,25$, $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2N\mathbb{Z}$ with $N=5,6$ $\mathbb{Z}/16\mathbb{Z}$
$p-1$ is not divisible by 333 $\mathbb{Z}/N\mathbb{Z}$ with $N=11,13,15,16,25$, $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2N\mathbb{Z}$ with $N=5,6,8$, $\mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}$ $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/16\mathbb{Z}$
Any prime $\mathbb{Z}/N\mathbb{Z}$ with $N=11,13,14,15,16,18,19,25,43,67,163$, $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/2N\mathbb{Z}$ with $N=5,6,7,8,9$, $\mathbb{Z}/5\mathbb{Z}\times\mathbb{Z}/5\mathbb{Z}$ $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/16\mathbb{Z}$
@@ 6. Realization of Torsion Subgroups
Let $E/\mathbb{Q}$ be an elliptic curve. Let $p>3$ be a prime such that $p-1$ is not divisible by 3,4,5,7,113,4,5,7,113 , 4 , 5 , 7 , 11. If $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is isomorphic to one of the following groups $\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/10\mathbb{Z},\mathbb{Z}/2\mathbb{Z}\times\mathbb{Z}/12\mathbb{Z}$ or $\mathbb{Z}/16\mathbb{Z}$, then $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}=E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}$.
@@
Assume that $E$ is in short Weierstrass form. Then $\mathbb{Q}(E[2])$ is the splitting field of the cubic $x^{3}+Ax+B$. Since $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ has full 222-torsion, the field $\mathbb{Q}(E[2])$ must be a subfield of $\mathbb{Q}(\zeta_{p})$. Then the cubic cannot be irreducible, since $p-1$ is not divisible by 333. This means that either the cubic has three rational roots, or it has one rational root and two other roots from $\mathbb{Q}(\sqrt{-p})$. This implies that $E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}$ also has full 222-torsion. Ultimately, $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}=E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}$.
Secondly, let $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}\cong\mathbb{Z}/16\mathbb{Z}$. Let $P=(x_{0},y_{0})\in E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ be a point of order 161616. Then, the extension degree of $\mathbb{Q}(x_{0})$ is equal to the number of elements in the set $\{x_{0}^{\sigma}:\sigma\in\operatorname{Gal}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})\}$, and it also divides $n$. In total, there are exactly 888 points of order 161616 in $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$, and their $x$-coordinates form a set of 444 elements. This means that $\{x_{0}^{\sigma}:\sigma\in\operatorname{Gal}(\mathbb{Q}(\zeta_{p})/\mathbb{Q})\}$ has at most 444 elements. It cannot have 333 or 444 elements because we assumed $n$ is not divisible by them. This implies that $\mathbb{Q}(x_{0})$ is at most a degree 222 extension, so $x_{0}\in\mathbb{Q}(\sqrt{-p})$. Also, $[\mathbb{Q}(x_{0},y_{0}):\mathbb{Q}(x_{0})]\leq 2$, so the extension degree of $\mathbb{Q}(x_{0},y_{0})$ is a divisor of 444, but it also divides $n$, so it is at most a degree 222 extension. This implies $P\in E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}$. Since $P$ generates $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$, we obtain $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}=E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}$. ∎
This theorem shows that if $E(\mathbb{Q}(\zeta_{p}))_{\text{tors}}$ is not one of the groups in Mazur’s theorem, then we should consider only the quadratic extension and look for torsion points there. These three groups appear in Theorem 2.2, so we know that they are realizable over infinitely many elliptic curves. However, we are interested in whether they are realizable over the quadratic fields $\mathbb{Q}(\sqrt{-p})$ for primes $p$ such that $p-1$ is not divisible by 333, 444, 555, 777, or 111111.
@@
[…]
The realization of $\mathbb{Z}/16\mathbb{Z}$ as a torsion subgroup over $\mathbb{Q}(\zeta_{p})$ remains an open problem in general. In order for this group to appear as a torsion subgroup, we need to find a prime $p$ such that $E(\mathbb{Q}(\sqrt{-p}))_{\text{tors}}\cong\mathbb{Z}/16\mathbb{Z}$ when $p\equiv 3\pmod{4}$. According to Theorem 1.3 in Banwait and Derickx [1], such a realization does not occur for any prime $p<800$ with $p\equiv 3\pmod{4}$. This rules out all the primes smaller than 800800800 considered in Theorem 1.1.
Their result further shows that among the primes $p\equiv 1\pmod{4}$, the only prime less than 800800800 for which this torsion structure is realized is $p=41$. For example, the elliptic curve 266910.ck6 satisfies $E(\mathbb{Q}(\sqrt{41}))_{\text{tors}}\cong\mathbb{Z}/16\mathbb{Z}$ (available on LMFDB). This fact, combined with Theorem 2.3, implies that the torsion subgroup $E(\mathbb{Q}(\zeta_{41}))_{\text{tors}}$ is isomorphic to $\mathbb{Z}/16\mathbb{Z}$.
Now, we turn our attention to the subgroups that arise when we relax the conditions in Theorem 1.1 and are realized over only finitely many elliptic curves. These groups are of the form $\mathbb{Z}/N\mathbb{Z}$, where the modular curve $X_{0}(N)$ has genus greater than zero and possesses non-cuspidal rational points. For example, we are not interested in the case $N=20$, even though $X_{0}(20)$ has genus one and six rational points, because all of these points are cusps. Therefore, there is no rational elliptic curve with a 202020-isogeny.
@@
where $A,B$ are rationals and $d$ is a square-free integer. This implies that the field of definition $\mathbb{Q}(P_{d})$ is contained in $\mathbb{Q}(P,\sqrt{d})$. As an example, we have the elliptic curve 121.a2 and its 111111-torsion is defined over $\mathbb{Q}(\zeta_{11})$. If we consider its 11-11- 11-quadratic twist 121.c1, then its 111111-torsion is defined over $\mathbb{Q}(\zeta_{11})^{+}$. If we consider its 3-3- 3-quadratic twist 1089.i2, then its 111111-torsion is defined over $\mathbb{Q}(\zeta_{33})^{+}$. This data is available on LMFDB.
@@
[…]
@@
If $j(E)=-7\cdot 137^{3}\cdot 2083^{3}$, the situation becomes more intriguing. We know that the field of definition of the 373737-torsion is an abelian number field contained in $\mathbb{Q}(\zeta_{n})$ for some positive integer $n$. To determine this minimal value of $n$, we apply Corollary 4.6, using the conductor $N_{E}=5^{2}\cdot 7^{2}$ of the minimal quadratic twist available on the LMFDB. The database also indicates that the extension degree of the field of definition is 363636 over $\mathbb{Q}$. From this, we deduce that $n$ must divide 129512951295.
To confirm that $n$ cannot be smaller, we use SageMath to factor the 373737-division polynomial of $E$ over $\mathbb{Q}$. This polynomial factors into four irreducible components over $\mathbb{Q}$: one of degree 181818 and three of degree 222222222. Let $P=(x_{0},y_{0})\in E(\overline{\mathbb{Q}})_{\text{tors}}$ be the generator of the kernel of the 373737-isogeny from $E$ to another elliptic curve. Then the degree $[\mathbb{Q}(x_{0}):\mathbb{Q}]$ must divide 363636. Hence, $\mathbb{Q}(x_{0})$ has degree 181818, and its minimal polynomial corresponds to the first factor of the 373737-division polynomial.
We then examine the splitting field of this polynomial and observe that 555, 777, and 373737 are ramified in the extension. This confirms that the minimal $n$ for which the field of definition is contained in $\mathbb{Q}(\zeta_{n})$ is indeed 129512951295 for the minimal quadratic twist corresponding to this $j$-invariant.
@@
14 $-3^{3}\cdot 5^{3}$ 7-7- 7 49.a2 $7^{2}$ $\mathbb{Q}(\zeta_{7})^{+}$
$3^{3}\cdot 5^{3}\cdot 17^{3}$ 7-7- 7 49.a1 $7^{2}$ $\mathbb{Q}(\zeta_{7})^{+}$
@@
$3^{3}\cdot 5^{3}/2$ Minimal 162.b4 $2\cdot 3^{4}$ $\mathbb{Q}(\zeta_{9})$
@@
We have already discussed $\mathbb{Z}/16\mathbb{Z}$ earlier in this chapter and provided an example where 161616-torsion is realized over $\mathbb{Q}(\zeta_{41})$. For the other cases, the elliptic curve 147.c2 satisfies $E(\mathbb{Q}(\zeta_{7})^{+})_{\text{tors}}\cong\mathbb{Z}/13\mathbb{Z}$, the elliptic curve 14.a4 satisfies $E(\mathbb{Q}(\zeta_{7})^{+})_{\text{tors}}\cong\mathbb{Z}/18\mathbb{Z}$, and the elliptic curve 11.a3 satisfies $E(\mathbb{Q}(\zeta_{11})^{+})_{\text{tors}}\cong\mathbb{Z}/25\mathbb{Z}$.